QS 1-8 Applying the accounting equation LO A1 1. Use the accounting equation to compute the missing financial statement amounts. pany Assets LiabilitiesEquity S 92,000 S 48,500 S 33,500 87,000 $ 119,000S 37,000 2. Use the expanded accounting equation to compute the missing financial statement amounts Revenues Expenses Liabilities Owner, Capital $ 24,500$ 37,000$ Assets 16,500 $ 74,000 $41,000 26,500 2 S 112,300S4 40,500s 61,000
Answer #1

1. Company 1 : Liabilities= 92000-48500= 43,500

  Company 2: Assets= 33500+87000= 120,500

Company 3: Equity= 119000-37000= 82,000

2. Company 1: Revenues= 74000-24500-37000+ 16500= 29,000

Company 2: Owners withdrawals= 40500+61000+41000-26500-112300=3,700